OpenStudy (anonymous):

If the annual interest rate is 10%, what is the approximate monthly interest rate

OpenStudy (anonymous):

\[\sqrt[12]{10\%}\] or equivalently \[(10\%)^\frac{1}{12}\]

OpenStudy (vincent-lyon.fr):

The idea is correct, but the way to compute is not. You need to start with \(\sqrt[12]{110\%}\)